The Subject: BOND VALUATION

The Nesmith Corporation's outstanding bonds have a $1,000 par value, a 12% semiannual coupon, 16 years to maturity, and an 8% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Solving problem with manual formula and calculator both:

Using financial calculator BA II Plus - Input details:

#

I/Y = R = Rate or yield / frequency of coupon in a year = 8/2 =

                    4.000000

PMT = Coupon rate x FV / frequency = -1000 x 12%/2 =

-$60.00

N = Number of years remaining x frequency = 16 x 2 =

32.00

FV = Future Value =

-$1,000.00

CPT > PV = Present value of bond = Price of Bond = Current value of bond =

$1,357.47

Formula for bond value: PV = |PMT| x ((1-((1+R%)^-N)) / R%) + (|FV|/(1+R%)^N) =

PV = (60* ((1-(1+0.04)^-32)/0.04) + 1000/(1+0.04)^32) =

$1,357.47